OpenStudy (anonymous):

Evaluate P(6, 2). 12 30 360

OpenStudy (anonymous):

Is that the full question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is it about polynomials ?

OpenStudy (anonymous):

no Permutations

OpenStudy (anonymous):

Do you know what factorials are ?

OpenStudy (anonymous):

what

OpenStudy (ryan123345):

GO @Eyad !!!

OpenStudy (anonymous):

@Shannon20150 he meant what is meant by factorial of a number. I mean, what would you understand by 4 factorial

OpenStudy (anonymous):

\[\Large P(6,2) = {6! \over (6 - 2)!}\] Is it like that ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Eyad yeah i think its that.

OpenStudy (anonymous):

30

OpenStudy (anonymous):

you have six items and you want to know how many ways you can order two of them counting principle tells you it is \(6\times 5\) i.e. six choices for the first one and then five for the second

OpenStudy (anonymous):

oh thanks @satellite73

OpenStudy (anonymous):

\[\Huge \color{Red}{OR}\] \[\Large P(6,2) = {6! \over (6 - 2)!} \] \[\Large P(6,2) = {6*(6-1)! \over (6 - 2)!}\] \[\Large P(6,2) = {6*(6-1)*(6-2)! \over (6 - 2)!}\] Cancel the (6-2)! 6*(6-1)=30 ---- Cheer !